# On-chip microwave coherent source with in-situ control of the photon number distribution

### v2

@@ II.2 Cavity control through tunable coupling
The model described above reveals that the cavity population is strongly dependent on the specific choices of the parameters $\mathrm{\Gamma}$, $\mathrm{\kappa}$, and $\mathrm{g}$ in a masing system. Consequently, the tunability of one of these parameters enables a complete control of the steady-state photon population in the target cavity, within a range that depends on the tunability of one or more of these parameters and on the others. In this study, we explore the implementation of a tunable coupler scheme [32] to manipulate the coupling strength $\mathrm{g}$ between the photon source and a target reservoir. The proposed configuration, depicted in Fig. 1(a-b), comprises a two-level system coupled to a superconducting cavity through a flux-tunable qubit. We can effectively model the dynamics of this system by using a Hamiltonian that incorporates the free evolution terms of its constituent elements and the exchange interactions until the next nearest neighbors. Taking into account just the two levels in the photon source spectra forming the transition protagonist of the resonant interaction with the target cavity, the Hamiltonian takes the following form:
$\mathrm{\frac{\hat{H}}{\hbar}=\hat{H}_{s}+\hat{H}_{c}+\hat{H}_{r}+\hat{H}_{rc}+\hat{H}_{sc}=}$ (4)
$\mathrm{\sum_{i=s,c}\frac{\omega_{i}\hat{\sigma}_{i}^{z}}{2}+\omega_{s}\hat{a}^{\dagger}\hat{a}+g_{rc}(\hat{a}^{\dagger}\hat{\sigma}_{c}^{-}+\hat{a}\hat{\sigma}_{c}^{+})+}$
$\mathrm{+g_{sc}(\hat{\sigma}_{s}^{+}\hat{\sigma}_{c}^{-}+\hat{\sigma}_{s}^{-}\hat{\sigma}_{r}^{+})+g_{sr}(\hat{\sigma}_{s}^{+}\hat{\sigma}_{c}^{-}+\hat{\sigma}_{s}^{-}\hat{\sigma}_{r}^{+})}\,.$
@@
Here, $\mathrm{\omega_{j}}$, $\mathrm{\hat{\sigma}_{j}^{z}}$, $\mathrm{\hat{\sigma}_{j}^{+}}$, and $\mathrm{\hat{\sigma}_{j}^{-}}$ for $\mathrm{j=s,c}$, represent the transition frequency and the Z-Pauli, raising and lowering operators for the photon source and the tunable coupler, respectively. Additionally, $\mathrm{\hat{a}}$ and $\mathrm{\hat{a}^{\dagger}}$ are the ladder operators associated with the target resonator modes. The terms $\mathrm{g_{ij}}$ for $\mathrm{i\neq j=s,r,c}$ represent the coupling strengths governing the interactions between the different elements. Dynamic control over the coupling between the photon source and the cavity can be achieved by engineering a dispersive coupling with the tunable coupler for both the photon source and the cavity, where $\mathrm{g_{jc}\ll\Delta_{jc}=\omega_{j}-\omega_{c}}$ for $\mathrm{j=s,r}$. Under this assumption, we can transform the Hamiltonian in Eq. 4 using a Schrieffer-Wolff transformation through the unitary operator $\mathrm{\hat{U}=\exp\left[\frac{g_{sc}}{\Delta_{sc}}(\sigma_{s}^{+}\sigma_{c}^{-}-\sigma_{s}^{-}\sigma_{c}^{+})+\frac{g_{rc}}{\Delta_{rc}}(a^{\dagger}\sigma_{c}^{-}-a\sigma_{c}^{+})\right]}$ that decouples the coupler from the other elements up to second order in $\mathrm{g/\Delta}$. This transformation results in a Jaynes-Cummings Hamiltonian with a flux-tunable coupling strength (further details in the Appendix)
$\mathrm{\frac{\hat{U}\hat{H}\hat{U}^{\dagger}}{\hbar}=\frac{\tilde{\omega}_{s}\hat{\sigma}_{s}^{z}}{2}+\tilde{\omega}_{r}\hat{a}^{\dagger}a+g_{eff}(\hat{\sigma}_{s}^{+}\hat{a}+\hat{\sigma}_{s}^{-}\hat{a})}\,.$ (5)
In this context, the influence of the capacitive couplings to the tunable coupler is quantitatively expressed by the renormalization of the characteristic frequencies governing the dynamics of the system. Specifically, $\mathrm{\tilde{\omega}_{k}=\omega_{k}+\frac{g_{kc}^{2}}{\Delta_{kc}}}$ with $\mathrm{k=s,r}$ represent the Lamb-shifted qubit and reservoir frequency, while $\mathrm{g_{eff}=g_{sc}g_{rc}\left(1/\Delta_{sc}+1/\Delta_{rc}\right)}$ characterizes the effective coupling between the photon source and the reservoir. The dependence on $\mathrm{\omega_{c}}$ embedded within $\Delta_{sc}$ and $\mathrm{\Delta_{rc}}$ offers a knob to regulate the coupling strength $\mathrm{g_{eff}}$ between the source and the cavity through an external flux modulating the coupler frequency. The result is an intrinsically quantum photon source at cryogenic temperatures, enabling on-demand control of photon flow into a target cavity through an external flux threading the coupler. The key factors for the practical feasibility of this scheme are the selection of the initial working point and the robustness against unwanted flux and design variations. The working point is determined by the choice of the counter-relaxation rate $\mathrm{\Gamma}$ and the loss rate of the target cavity $\mathrm{\kappa}$. As highlighted in Eq. 3, the selection of these two parameters dictates whether the system can coherently inject photons into the target cavity and also determines the maximum number of steady-state photons that can be stored within the target cavity in the masing regime. The degree of tunability with the external flux $\Phi_{c}$ over the steady-state photon distribution is evaluated by examining the second-order correlation function $\mathrm{g^{(2)}(0)=\braket{\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}}/\braket{\hat{a}^{\dagger}\hat{a}}}$. The latter is a commonly used metric for distinguishing quantum effects from classical phenomena in quantum optics systems [41]. In the context of this scheme, we use it as an indicator of the operational regime of the system [31], as illustrated in Fig. 1(c). When $\mathrm{g^{(2)}(0)=1}$, the system is in the masing regime, where the target cavity gets populated by $\mathrm{\Gamma/2\kappa}$ photons through its interaction with the photon source, while for $\mathrm{g^{(2)}(0)=2}$ the system is in the thermal regime where the masing regime is suppressed and there is no effective population of the cavity. The suppression of the masing regime can occur when the coupling between the photon source and the cavity decreases to a point where the loss rate of photons equals or exceeds the photon injection rate into the cavity. To maximize the range of photon distributions that can be generated in the cavity, the system must be engineered in a way that allows to pass from the masing to the thermal regime within the available range of effective coupling. Considering a generic working point $\mathrm{\left(\Gamma,\kappa\right)}$ with an effective coupling that can be tuned inside a certain range of values, the more the system can span between $\mathrm{g^{(2)}(0)=1}$ and $\mathrm{g^{(2)}(0)=2}$, the wider is the range of average steady-state photons that can be injected in the cavity. To quantitatively assess the working area of the system, we compute the second-order correlation function using the model in Eq. 5 with QuTip [53] and estimate the variation of $\mathrm{g^{(2)}(0)}$ within a certain range of effective coupling. Considering Eq. 3, we evaluate the variation of the second-order correlation function between the cases where $\mathrm{g_{eff}=\sqrt{\Gamma\kappa/8}}$ and $\mathrm{g_{eff}=\sqrt{\Gamma\kappa/0.4}}$ which, in good approximation, correspond to two situations where the cavity is maximally and minimally populated, respectively. The final results are depicted in Fig. 2, where the colormap highlights the expected variations of $\mathrm{g^{(2)}(0)}$ inside a range of experimentally reachable working points of the system [30, 50]. The red points in the colormap represent the maximum variation of $\mathrm{\Delta g^{(2)}}$ indicating the working points where the number of steady-state photons
Figure 2: Variation of the second-order correlation function $\mathrm{g^{(2)}(0)}$ as a function of the counter-relaxation rate $\mathrm{\Gamma}$ and the loss rate in the target cavity $\mathrm{\kappa}$, while considering a range of coupling strengths from $\mathrm{g_{eff}=\sqrt{\Gamma\kappa/8}}$ to $\mathrm{g_{eff}=\sqrt{\Gamma\kappa/0.4}}$. The white circle marks the working point used in this study for validating the photon source scheme. The role of the flux $\mathrm{\Phi_{c}}$ tuning the coupling $\mathrm{g_{eff}}$ in this point is figuratively represented as a flux-dependent knob, allowing the transition between deep Masing (M) and Thermal (T) regime. This control mechanism regulates consequently the average number of steady-state photons injected into the target reservoir.

### v3

@@ II.3 Control through tunable coupling
In this study, we explore the implementation of a tunable coupler scheme [46] to manipulate the coupling strength $\mathrm{g_{eff}}$ between the photon source and a target reservoir. The proposed configuration, depicted in Fig. 1(a-b), comprises a two-level system coupled to a superconducting reservoir through a flux-tunable qubit. By employing a broadband coaxial line inductively coupled to the SQUID in the tunable coupler, flux can be delivered on-chip, enabling control over the effective coupling strength between the photon source and the target reservoir without directly interfacing with either element. The circuit can also be integrated into a Single Flux Quantum (SFQ) architecture [68], in which flux is provided via dedicated voltage-biased superconducting circuits, removing the need for a broadband coaxial line. We model the dynamics of this system by using a Hamiltonian that incorporates the free evolution terms of its constituent elements and the exchange interactions until the next nearest neighbors. Taking into account just the first two levels of the tunable coupler, the Hamiltonian takes the following form:
$\hat{H}/\hbar$ $=\left(\hat{H}_{s}+\hat{H}_{c}+\hat{H}_{r}+\hat{H}_{rc}+\hat{H}_{sc}+\hat{H}_{sr}\right)/\hbar$ (4)
$=\sum_{i=s,c}\frac{\omega_{i}\hat{\sigma}_{i}^{z}}{2}+\omega_{r}\hat{a}^{\dagger}\hat{a}+g_{rc}(\hat{a}^{\dagger}\hat{\sigma}_{c}^{-}+\hat{a}\hat{\sigma}_{c}^{+})$
$\quad+g_{sc}(\hat{\sigma}_{s}^{+}\hat{\sigma}_{c}^{-}+\hat{\sigma}_{s}^{-}\hat{\sigma}_{r}^{+})+g_{sr}(\hat{\sigma}_{s}^{+}\hat{a}+\hat{\sigma}_{s}^{-}\hat{a}^{\dagger})\,.$
@@
Here, $\mathrm{\omega_{j}}$, $\mathrm{\hat{\sigma}_{j}^{z}}$, $\mathrm{\hat{\sigma}_{j}^{+}}$, and $\mathrm{\hat{\sigma}_{j}^{-}}$ for $\mathrm{j=s,c}$, represent the transition frequency and the Z-Pauli, raising and lowering operators for the photon source and the tunable coupler, respectively. Additionally, $\mathrm{\hat{a}}$ and $\mathrm{\hat{a}^{\dagger}}$ are the ladder operators associated with the target resonator modes. The terms $\mathrm{g_{ij}}$ for $\mathrm{i\neq j=s,r,c}$ represent the coupling strengths governing the interactions between different elements, defined in terms of the circuital parameters of the scheme:
$\mathrm{g_{ic}}$ $\approx\frac{1}{2}\frac{C_{ic}}{\sqrt{C_{i}C_{c}}}\sqrt{\omega_{i}\omega_{c}}\,,$ (5)
$\mathrm{g_{sr}}$ $\approx\frac{1}{2}\frac{C_{sc}C_{rc}}{\sqrt{C_{s}C_{r}C_{c}^{2}}}\sqrt{\omega_{s}\omega_{r}}\,,$
where $\mathrm{C_{ij}}$ for $\mathrm{i\neq j=s,r,c}$ correspond to the coupling capacitor between the different elements in the scheme, while $\mathrm{C_{i}}$ for $\mathrm{i=s,r,c}$ denote their capacitance to ground. The coupling strengths $\mathrm{g_{rc}}$ and $\mathrm{g_{sc}}$ associated with the interactions involving the tunable coupler can be controlled via the magnetic flux $\mathrm{\Phi_{c}}$, which modulates the frequency of the tunable coupler $\mathrm{\omega_{c}\left(\Phi_{c}\right)}$. In contrast, the coupling strength $\mathrm{g_{sr}}$ related to the next-nearest neighbor interaction between the source and the reservoir, remains fixed. Flux tunability of the coupling between the photon source and the reservoir can be realized by engineering a dispersive coupling between the tunable coupler and both the photon source and the reservoir when $\mathrm{g_{jc}\ll\Delta_{jc}=\omega_{j}-\omega_{c}}$ for $\mathrm{j=s,r}$. Under these assumptions, the Hamiltonian in Eq. 4 can be simplified using a Schrieffer-Wolff transformation through the unitary operator $\mathrm{\hat{U}=\exp\left[\sum_{j=s,r}\frac{g_{jc}}{\Delta_{jc}}(\hat{\sigma}_{j}^{+}\hat{\sigma}_{c}^{-}-\hat{\sigma}_{j}^{-}\hat{\sigma}_{c}^{+})\right]}$. This transformation effectively decouples the tunable coupler from the photon source and reservoir to the second order in $\mathrm{g_{sc}/\Delta_{sc}}$ and $\mathrm{g_{rc}/\Delta_{rc}}$. This transformation results in a Jaynes-Cummings Hamiltonian with a flux-tunable coupling strength (further details in Methods ”Hamiltonian”)
$\mathrm{\frac{\hat{U}\hat{H}\hat{U}^{\dagger}}{\hbar}=\frac{\tilde{\omega}_{s}\hat{\sigma}_{s}^{z}}{2}+\tilde{\omega}_{r}\hat{a}^{\dagger}\hat{a}+g_{eff}(\hat{\sigma}_{s}^{+}\hat{a}+\hat{\sigma}_{s}^{-}\hat{a})}\,.$ (6)
In this framework, the influence of the tunable coupler is quantitatively expressed by the renormalization of the characteristic frequencies governing the system dynamics. Specifically, $\mathrm{\tilde{\omega}_{k}=\omega_{k}+\frac{g_{kc}^{2}}{\Delta_{kc}}}$ with $\mathrm{k=s,r}$ represent the Lamb-shifted qubit and reservoir frequency, while the effective coupling $\mathrm{g_{eff}}$ between the source and the reservoir is defined by the following expression:
$\mathrm{g_{eff}(\Phi_{c})=g_{sc}g_{rc}\left(\frac{1}{\Delta_{sc}}+\frac{1}{\Delta_{rc}}\right)+g_{sr}}\,.$ (7)
The effective coupling $\mathrm{g_{eff}}$ is defined by the sum of two contributions arising from the interactions between the eigenstates |$\mathrm{\ket{s,c,r}}$ inside the system spectra. The first contribution comes from the virtual interaction between the first excited state of the coupler |$\mathrm{\ket{g,1,n_{ph}}}$ with the states |$\mathrm{\ket{e,0,n_{ph}}}$ and |$\mathrm{\ket{g,0,n_{ph}+1}}$ responsible for the photon transport from the source to the reservoir. When the source and the reservoir are on resonance, the eigenstates |$\mathrm{\ket{e,0,n_{ph}-1}}$ and |$\mathrm{\ket{g,0,n_{ph}}}$ are degenerate and form a pair of dressed states split by an amount $\mathrm{\delta_{N}=2g_{sr}\sqrt{N}}$. Similarly, the eigenstate with the coupler in the excited state |$\mathrm{\ket{g,1,n_{ph}}}$ forms a pair of dressed states with |$\mathrm{\ket{e,1,n_{ph}-1}}$. The splitting $\mathrm{\delta_{N}}$ becomes comparable to the minimal detuning between energy levels $\mathrm{\left(\Delta_{m}=min\left[\Delta_{rc},\Delta_{sc}\right]\right)}$ for $\mathrm{N^{*}=\left(\Delta_{min}/g_{sr}\right)^{2}}$. Within the dispersive regime, this number $\mathrm{N^{*}}$ always lies outside the range of photons that can be injected before the dispersive regime ceases to be valid. As a result, we can reasonably assume that the virtual interaction remains independent of the number of photons in the reservoir throughout the entire injection process (further details in Methods ”Functional dependence on the number of photons”). The second term accounts for the nearest neighbor coupling that directly couples the two states |$\mathrm{\ket{e,0,n_{ph}-1}}$ and |$\mathrm{\ket{g,0,n_{ph}}}$. The dependence on $\mathrm{\omega_{c}}$ embedded within $\mathrm{\Delta_{sc}}$ and $\mathrm{\Delta_{rc}}$ offers a knob to regulate the coupling strength $\mathrm{g_{eff}}$
Figure 2: Microwave emission tunability via tunable coupling The average number of steady-state photons $\mathrm{n_{ph}}$ and the normalized second-order correlation function $\mathrm{g^{(2)}(0)}$ are shown in terms of the effective coupling strength $g_{\text{eff}}$ between the photon source and the reservoir and in terms of the masing ratio $\Gamma\kappa/4g_{eff}^{2}$ where $\kappa$ is the loss rate of the reservoir. The results are obtained by simulating the steady-state dynamics of the Hamiltonian in Eq. 6 in Qutip varying the effective coupling $\mathrm{g_{eff}}$. The dimension $\mathrm{N}$ of the Hilbert subspace associated with the harmonic reservoir in the simulation is $\mathrm{N=100}$.
between the source and the reservoir through an external flux modulating the coupler frequency. The result is an intrinsically quantum photon source at cryogenic temperatures, enabling on-demand control of photon flow into a target cavity through an external flux threading the coupler.
The practical feasibility of the proposed scheme relies on several key factors. We address them by examining the different engineering steps required for its implementation. The first step is the selection of the counter-relaxation rate $\mathrm{\Gamma}$ and the loss rate of the target reservoir $\mathrm{\kappa}$ which together define the working point $\mathrm{\left(\Gamma,\kappa\right)}$ of the scheme. According to Eq. 3, these parameters determine whether the system can coherently inject photons for a given value of the effective coupling $\mathrm{g_{eff}}$, as well as the maximum emission, which equals to $\mathrm{n_{ph}^{max}=\Gamma/2\kappa}$. Assuming that the flux-controlled effective coupling allows the masing ratio to vary within a certain range $\mathrm{\left[\lambda_{min},\lambda_{max}\right]}$, we evaluate the resulting tunability of the emission power, for given $\mathrm{\lambda_{min},\lambda_{max}}$. We perform this analysis by examining the normalized second-order correlation function $\mathrm{g^{(2)}(0)=\braket{\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}}/\braket{\hat{a}^{\dagger}\hat{a}}^{2}}$ [69], a common metric for distinguishing quantum effects from classical phenomena in quantum optical systems [69, 70]. Within this scheme, $\mathrm{g^{(2)}(0)}$ serves as an indicator of the system operational regime [67], as illustrated in Fig. 2. When $\mathrm{g^{(2)}(0)=1}$, the photon statistics follow a Poissonian distribution [69] and the emissions reach a maximum $\mathrm{\left(n_{ph}\sim\Gamma/2\kappa\right)}$, which denotes that the system is in the masing regime. Reducing the coupling between the photon source and the reservoir can suppress the masing dynamics, bringing the photon loss rate closer to the emissions rate. In this intermediate regime, the second-order correlation function assumes values between $\mathrm{1<g^{(2)}(0)<2}$ and the photon distribution follows a super-Poissonian statistics (more details in Methods ”From Poissonian to super-Poissonian”). When the losses dominate over the emissions and the masing is fully suppressed, the system is in a thermal regime $\mathrm{g^{(2)}(0)=2}$ [69] where the reservoir remains unpopulated [67].
[…]
Even when the system is well designed to operate at an optimal working point, it is essential to ensure robustness against fabrication errors, which may affect the range of emissions, and against flux noise, which can degrade control over the steady-state photon distribution. To evaluate the impact of design errors on system performance, we consider the previous reference working point $\mathrm{\left(\Gamma,\kappa\right)}$ and introduce errors of up to ±$\mathrm{\pm 20\%}$ in both parameters. In each case, we compute the variation $\mathrm{\Delta n_{ph}}$, defined as the difference in emission between the cases where the effective coupling $\mathrm{g_{eff}}$ is set to its maximum and minimum values. The maximum and minimum effective coupling values are given by $\mathrm{g_{eff}=\sqrt{\Gamma\kappa/0.4}}$ and $\mathrm{g_{eff}=\sqrt{\Gamma\kappa/8}}$, respectively, assuming the coupling is designed for the reference working point $\mathrm{\left(\Gamma,\kappa\right)}$. The results are depicted in Fig. 3(b) and show a fair robustness of the system within $\mathrm{5\%}$ of error in both degrees of freedom. To prevent the influence of faulty fabrication on the system feasibility, the coupling can be engineered to work within a broader range. This expansion of the coupling range, however, comes at the expense of an increased sensitivity of the tunable coupler to external flux noise.
In a practical realization of the system, a trade-off must be found between these two features. To decrease the sensitivity to external flux, we can employ several strategies that allow us to precisely define the range of working points. Differently from the typical tunable coupler scheme in which the system is engineered to have an off-point value where the coupling is cancelled [46], here we just aim at modulating it within a specific range of couplings between a deep masing regime $\mathrm{\Gamma\kappa/4g_{eff}^{2}\ll 1}$ to a suppressed masing regime $\mathrm{\Gamma\kappa/4g_{eff}^{2}>1}$. One approach involves designing the tunable coupler with a resonance frequency lower than that of the photon source and the reservoir, resulting in an effective coupling strength $\mathrm{g_{eff}\propto 1/(\omega_{s}-\omega_{c})}$. Additionally, we can use an asymmetric SQUID as the tunable element of the coupler, which exhibits a more gradual response to external fluxes [72]. These adjustments provide two significant advantages: (1) the coupler never approaches a frequency close to those of the photon source and the reservoir for any threaded
[…]
flux, ensuring that the dispersive condition is steadily met; (2) the asymmetry allows for non-zero minima in the coupler frequency, narrowing the range of effective coupling to well-defined maxima and minima.